OpenStudy (anonymous):

Linda has two fish tanks in the shape of rectangular prisms. Each of the sides of the second fish tank is 3 times greater than the sides of the first fish tank. If Linda fills both the fish tanks completely with water, the volume of water in the second fish tank will be

OpenStudy (anonymous):

This is a question asking for you to find the ratio of the volumes. To do so, you will need to know that the volume of a rectangular sphere is V=b*h*w. for the first fish tank (the smaller one), you just have V=b*h*w For the second fish tank (the one that is scaled by a factor of 3), you get V=3b*3h*3w. If we simplify this equation, we get V=3*3*3*b*h*w, or V=27*b*h*w comparing the volume equation to the equation for the first, smaller fish tank, we see that the volume is scaled by 27 times the first, or cubed the scale, 27 to 1, or 27/1 or whatever term you're familiar with. HTH!